OpenStudy (anonymous):

X^2+4X=5 COMPLETE THE SQUARE

OpenStudy (caozeyuan):

X^2+4X+4=7, so (X+2)^2=7

OpenStudy (anonymous):

\[x^2+4x=5\\ x^2+4x+4=5+4\\ (x+2)^2=9\]

OpenStudy (caozeyuan):

@SithsAndGiggles , sorry my typo, yes, the cons. is 9, not 7